# Shijian Li1,2, Tao Li4,5,1, Gang Pan1,2,5 1

### v1

@@ 1 Introduction
Brain-Computer Interface (BCI) refers to a system that allows direct communication between the brain and external devices or computers. Electroencephalography (EEG) is a technique used to measure and record electrical activity in the brain, where electrodes are placed on the scalp to detect and amplify the brain’s electrical signals. EEG plays a crucial role in BCI as it provides a non-invasive and real-time measure of brain activity that can be used to decode and interpret user’s intentions or commands. By analyzing the patterns and features in EEG signals, algorithms and models have been developed to decode specific brain states, including but not limited to emotion recognition (Dadebayev et al., 2022; Gao et al., 2024), motor imagery classification (Altaheri et al., 2021; Dai et al., 2020), seizure detection (Ahmad et al., 2022; Yıldız et al., 2022) and sleep staging (Phan & Mikkelsen, 2022; Wang et al., 2024).
@@
Impressed by self-supervised learning (SSL) (Liu et al., 2021c) on computer vision (CV) and natural language processing (NLP), some studies (Kostas et al., 2021; Chien et al., 2022; Yang et al., 2023; Foumani et al., 2024; Jiang et al., 2024) propose EEG foundation models which are pre-trained on a vast amount of EEG data with self-supervised learning and fine-tuned on downstream datasets for some clinical or BCI applications. Most of these methods segment the original EEG signals into patches of EEG channels to ensure that the neural network can deal with EEG signals with diverse channels and time length, shown in Figure 1(a). Consistent with the approach of handling image patches used in ViT (Dosovitskiy et al., 2020), they flatten the EEG patches and feed them into transformers to model the dependencies among all EEG patches, called full EEG modeling strategy shown in Figure 1(b). Meantime, they adopted absolute positional encoding based on electrode numbering as a channel embedding to encode positional information. However, existing methods ignore two characteristics of EEG signals: Unique structural characteristics: EEG signals have different structural characteristics compared to image. It contain heterogeneous spatial and temporal dependencies, but images contain only spatial dependencies. And the dependencies among EEG patches within the same channel or time interval could be stronger than those among patches from different channels and time intervals, but such a prior assumption may not be valid in image. The full EEG modeling strategy models the dependencies among all EEG patches together, ignoring the unique structural characteristics of EEG signals. Channel variation: EEG channels are not solely defined by electrode position but are also influenced by the referencing scheme used (e.g., earlobe, average, REST, or bipolar references). Existing EEG foundation models, such as LaBraM (Jiang et al., 2024), employ absolute positional encoding based on electrode numbering as a channel embedding, but this method assumes a fixed relationship between EEG channels and electrode positions. As a result, absolute positional encodings tied directly to electrode numbering may limit the model’s adaptability across tasks and datasets that vary in spatial and reference properties.
@@ 3.1 Pre-training
Pre-training Settings We implemented CBraMod based on the Python 3.11.7 and PyTorch 2.1.2 + CUDA 12.1. We set the time duration of each EEG patch as 1 second (200 data points), and one 30-second EEG sample is segmented to $19\times 30=570$ EEG patches. For the model configurations, the patch encoder consists of 3-layer 1D convolution with group normalization and GELU activation. The positional encoder is an one-layer 2D depthwise CNN. The backbone of CBraMod is a 12-layer criss-cross transformer with 200 hidden dimensions, 800 inner dimensions (feed-forward), and 8-head criss-cross attention (4 heads for S-Attention and 4 heads for V-Attention). 50% of mask ratio is used to randomly mask the patches. The batch size was set to 128 and the number of epochs was set to 40. The model is trained using the AdamW optimizer with default settings, the learning rate is set to 5e-4 and the weight decay is set to 5e-2. CosineAnnealingLR was used to dynamically adjust the learning rate during the pre-training. CBraMod was pre-trained on one machine with Intel Xeon Gold 6226R CPU and four NVIDIA RTX A5000 GPU for about 5 days. The source code and pre-trained weights will be publicly available. More details can be found in Appendix B.
@@ 3.2 Experiment Setup of Downstream BCI Tasks
Downstream BCI Tasks and Datasets To comprehensively evaluate the performance of our method, we select up to 10 downstream BCI tasks as follows: I. Emotion Recognition refers to detecting and interpreting emotional states or patterns based on EEG. We choose FACED (Chen et al., 2023) and SEED-V (Liu et al., 2021b) as downstream datasets for this task. II. Motor Imagery Classification is a process that involves decoding or classifying motor imagery tasks based on brain activity patterns captured through EEG. We use PhysioNet-MI (Schalk et al., 2004) and SHU-MI (Goldberger et al., 2000b) to evaluate the performance of CBraMod on this task. III. Sleep Staging, also known as sleep stage classification, categorizes different stages of sleep based on physiological signals (e.g. EEG) recorded during sleep. ISRUC (Khalighi et al., 2016) is a commonly used benchmark for sleep staging. IV. Seizure Detection refers to the process of identifying and recognizing epileptic seizures based on physiological signals, such as EEG. Here, CHB-MIT (Shoeb, 2009) is choosed for this task. V. Imagined Speech Classification involves the identification and interpretation of imagined speech or covert speech intentions based on brain activity patterns such as EEG. BCIC2020-3 (Jeong et al., 2022) is a dataset for this task on 2020 international brain–computer interface competition. VI. Mental Disorder Diagnosis refers to the process of identifying and categorizing mental health states based on EEG. Mumtaz2016 (Mumtaz, 2016) is an EEG dataset collected from patients with major depressive disorder and normal controls. VII. Vigilance Estimation refers to the assessment and measurement of an individual’s level of vigilance or sustained attention from EEG. SEED-VIG (Min et al., 2017) is dataset used to assess the level of vigilance. VIII. Mental Stress Detection involves using EEG to identify an individual’s level of stress. MentalArithmetic (Zyma et al., 2019) is a dataset containing EEG recordings of subjects before and during the performance of mental arithmetic tasks. IX. EEG Events Classification involves the categorization and identification of specific events or patterns in EEG. We choose TUEV (Obeid & Picone, 2016) for evaluation, which was consistent with BIOT (Yang et al., 2023) and LaBraM (Jiang et al., 2024). X. Abnormal Detection refers to the process of identifying and detecting abnormal patterns or events in EEG. Similarly, TUAB (Obeid & Picone, 2016) is used for evaluation, consistent with BIOT (Yang et al., 2023) and LaBraM (Jiang et al., 2024). The corresponding datasets of all downstream BCI tasks are shown in Table 1. For all the downstream datasets, we resampled the EEG signals into 200 Hz and set the time duration of each EEG patch as 1 second (200 data points), consistent with the pre-training data. More details of each dataset and its preprocessing are introduced in Section 3.3 and Appendix E.
@@ Appendix A Related Work
EEG Decoding. EEG is a non-invasive technique to measure brain activity. Early studies on EEG decoding predominantly employed traditional machine learning methods (Bashashati et al., 2007; McFarland et al., 2006; Lotte et al., 2007), which usually depend on hand-crafted features that require lots of prior knowledge and could have weak generalizability. With the development of deep learning (LeCun et al., 2015) techniques, an increasing number of researchers are shifting their focus to studying EEG decoding methods based on deep learning (Parvaneh et al., 2019; Craik et al., 2019; Al-Saegh et al., 2021; Sekkal et al., 2022; Yang et al., 2022). For example, some studies utilize convolutional neural network (CNN) to extract temporal and spatial features from EEG for different BCI tasks like motor imagery classification, emotion recognition and seizure detection (Schirrmeister et al., 2017; Sakhavi et al., 2018; Lawhern et al., 2018; Abdelhameed & Bayoumi, 2021; Ding et al., 2022). Long Short-Term Memory (LSTM) is also used for EEG feature extraction and classification on BCI tasks such as motor imagery classification and sleep staging (Wang et al., 2018; Phan et al., 2019). Some researchers propose CNN-LSTM to learn EEG features for motor imagery classification, emotion recognition and sleep staging (Supratak et al., 2017; Zhang et al., 2019; Dar et al., 2020; Li et al., 2022; Wang et al., 2023a), where CNN is usually used to extract local features and LSTM is utilized to capture global dependencies. Transformer architecture is also utilized to learn spatial-temporal feature for BCI tasks including emotion recognition, sleep staging and person identification (Song et al., 2021; Liu et al., 2021a; Du et al., 2022; Phan et al., 2022). To combine the strengths of CNN and transformer, some works devise a CNN-Transformer network for automated artifact detection and other tasks (Song et al., 2022; Peh et al., 2022; Wang et al., 2023b; Zhou et al., 2024). In addition, some studies use graph neural networks learn spatial-temporal features from multi-channel EEG for various BCI tasks (Jia et al., 2020; 2021; Ding et al., 2023). These methods perform well on specific tasks or datasets, but collecting labeled EEG data is costly and labor-intensive. Meantime, the variations in EEG signal formats across different datasets pose challenges for enhancing model performance and generalizability.
@@
Brain Foundation Model. A foundation model (Bommasani et al., 2021) is a deep learning model trained on extensive data, typically using self-supervision at a large scale, and can be adapted, such as through fine-tuning, for various downstream tasks. Foundation models, such as BERT (Devlin et al., 2018), MAE (He et al., 2022), CLIP (Radford et al., 2021), GPT-4 (Achiam et al., 2023), SAM (Kirillov et al., 2023) and SORA (Brooks et al., 2024), have achieved remarkable success in computer vision, natural language, and multimodal, but there is still a vast unexplored potential for foundation models in brain signals. Some studies propose time series pre-trained models (Eldele et al., 2021; Zhang et al., 2022; Woo et al., 2024; Chen et al., 2024) based on contrastive learning or others for many scenarios (e.g., weather, traffic flow, exchange rates), one of which is brain signals. Some studies learn unsupervised representations from EEG by utilizing self-supervised learning pretext tasks to uncover the structure of clinical EEG signals (Banville et al., 2021). BENDER (Kostas et al., 2021) is proposed to address the problem of limited labeled data on EEG by using a contrastive self-supervised learning task to learn generic EEG representations. MAEEG (Chien et al., 2022) is masked auto-encoder for learning EEG representations by reconstruction-based self-supervised learning. BrainBERT (Wang et al., 2022) is a pre-training model for intracranial recordings, which learns a complex non-linear transformation of neural data by construction of the masked stereo-electroencephalographic (SEEG) spectrogram. Brant (Zhang et al., 2023) is a foundation model for intracranial neural signals which attends long-term dependency and captures spatial correlation across channels. BIOT (Yang et al., 2023) is a generic biosignal learning model which enables joint pre-training and knowledge transfer across different biosignal datasets in the wild. EEG2Rep (Foumani et al., 2024) is a self-prediction approach for self-supervised representation learning from EEG, which predicting masked inputs in the latent representation space. LaBraM (Jiang et al., 2024) is a large brain model, which learns EEG generic representations by predicting the corresponding neural tokens of masked EEG patches. However, most of existing EEG foundational model the spatial and temporal dependencies between all EEG patches together, ignoring the unique structural characteristics of EEG signals. Meanwhile, existing EEG foundation models have limited generalizability, performing well only on a limited range of downstream tasks.
@@ Appendix D More Details for Experimental Settings on Downstream BCI Tasks
D.1 Baselines
@@ E.9 Motor Imagery Classification
To further validate the performance of CBraMod on the motor imagery tasks, we conducted additional experiments on the BCIC-IV-2a (Brunner et al., 2008) dataset. The BCI Competition IV Dataset 2a, provided by Graz University of Technology, comprises EEG recordings from 9 subjects performing 4 motor imagery tasks: imagining movements of the left hand, right hand, both feet, and tongue. Data were collected over two sessions on separate days using 22 Ag/AgCl electrodes at a sampling rate of 250 Hz. Each session consisted of 288 EEG trials, with 72 trials per task. We used [2, 6] seconds of each trial. A band-pass filtered (0.3 Hz–40 Hz) was applied to remove the low-frequency and high-frequency noise. We resample the EEG signals to 200 Hz and obtain 5,088 4-second samples. We use a strict subject-indenpendent train/validation/test strategy as other MI datasets in our paper. Subject 1-5, 6-7, 8-9 are used for training, validation, and test, respectively. The results are shown in Table 15. CBraMod continues to outperform existing methods on this dataset, further reinforcing the generalizability of our method.

### v2

@@ 1 Introduction
Brain-Computer Interface (BCI) refers to a system that allows direct communication between the brain and external devices or computers. Electroencephalography (EEG) is a technique used to measure and record electrical activity in the brain, where electrodes are placed on the scalp to detect and amplify the brain’s electrical signals. EEG plays a crucial role in BCI as it provides a non-invasive and real-time measure of brain activity that can be used to decode and interpret user’s intentions or commands. By analyzing the patterns and features in EEG signals, algorithms and models have been developed to decode specific brain states, including but not limited to emotion recognition (Dadebayev et al., 2022; Gao et al., 2024), motor imagery classification (Altaheri et al., 2021; Dai et al., 2020), seizure detection (Ahmad et al., 2022; Yıldız et al., 2022) and sleep staging (Phan & Mikkelsen, 2022; Wang et al., 2024b; Zhou et al., 2024a).
@@
Impressed by self-supervised learning (SSL) (Liu et al., 2021c) on computer vision (CV) and natural language processing (NLP), some studies (Kostas et al., 2021; Chien et al., 2022; Yang et al., 2023; Foumani et al., 2024; Jiang et al., 2024; Wang et al., 2024a) propose EEG foundation models which are pre-trained on a vast amount of EEG data with self-supervised learning and fine-tuned on downstream datasets for some clinical or BCI applications. Most of these methods segment the original EEG signals into patches of EEG channels to ensure that the neural network can deal with EEG signals with diverse channels and time length, shown in Figure 1(a). Consistent with the approach of handling image patches used in ViT (Dosovitskiy et al., 2020), they flatten the EEG patches and feed them into transformers to model the dependencies among all EEG patches, called full EEG modeling strategy shown in Figure 1(b). Meantime, they adopted absolute positional encoding based on electrode numbering as a channel embedding to encode positional information. However, existing methods ignore two characteristics of EEG signals: Unique structural characteristics: EEG signals have different structural characteristics compared to image. It contain heterogeneous spatial and temporal dependencies, but images contain only spatial dependencies. And the dependencies among EEG patches within the same channel or time interval could be stronger than those among patches from different channels and time intervals, but such a prior assumption may not be valid in image. The full EEG modeling strategy models the dependencies among all EEG patches together, ignoring the unique structural characteristics of EEG signals. Channel variation: EEG channels are not solely defined by electrode position but are also influenced by the referencing scheme used (e.g., earlobe, average, REST, or bipolar references). Existing EEG foundation models, such as LaBraM (Jiang et al., 2024), employ absolute positional encoding based on electrode numbering as a channel embedding, but this method assumes a fixed relationship between EEG channels and electrode positions. As a result, absolute positional encodings tied directly to electrode numbering may limit the model’s adaptability across tasks and datasets that vary in spatial and reference properties.
@@ 3.1 Pre-training
Pre-training Settings We implemented CBraMod based on the Python 3.11.7 and PyTorch 2.1.2 + CUDA 12.1. We set the time duration of each EEG patch as 1 second (200 data points), and one 30-second EEG sample is segmented to $19\times 30=570$ EEG patches. For the model configurations, the patch encoder consists of 3-layer 1D convolution with group normalization and GELU activation. The positional encoder is an one-layer 2D depthwise CNN. The backbone of CBraMod is a 12-layer criss-cross transformer with 200 hidden dimensions, 800 inner dimensions (feed-forward), and 8-head criss-cross attention (4 heads for S-Attention and 4 heads for V-Attention). 50% of mask ratio is used to randomly mask the patches. The batch size was set to 128 and the number of epochs was set to 40. The model is trained using the AdamW optimizer with default settings, the learning rate is set to 5e-4 and the weight decay is set to 5e-2. CosineAnnealingLR was used to dynamically adjust the learning rate during the pre-training. CBraMod was pre-trained on one machine with Intel Xeon Gold 6226R CPU and four NVIDIA RTX A5000 GPU for about 5 days. More details can be found in Appendix B.
@@ 3.2 Experiment Setup of Downstream BCI Tasks
Downstream BCI Tasks and Datasets To comprehensively evaluate the performance of our method, we select up to 10 downstream BCI tasks. All the downstream BCI tasks with the corresponding datasets are presented in Table 1 and Appendix D.1. For all the downstream datasets, we resampled the EEG signals into 200 Hz and set the time duration of each EEG patch as 1 second (200 data points), consistent with the pre-training data. More details of each dataset and its preprocessing are introduced in Section 3.3 and Appendix E.
@@ 4 Conclusion
Acknowledgments
This work was supported by STI 2030 Major Projects (2021ZD0200400), Natural Science Foundation of China (No. 61925603) and the Key Program of the Natural Science Foundation of Zhejiang Province, China (No. LZ24F020004). The corresponding authors are Dr. Sha Zhao and Dr. Gang Pan.
@@ Appendix A Related Work
EEG Decoding. EEG is a non-invasive technique to measure brain activity. Early studies on EEG decoding predominantly employed traditional machine learning methods (Bashashati et al., 2007; McFarland et al., 2006; Lotte et al., 2007), which usually depend on hand-crafted features that require lots of prior knowledge and could have weak generalizability. With the development of deep learning (LeCun et al., 2015) techniques, an increasing number of researchers are shifting their focus to studying EEG decoding methods based on deep learning (Parvaneh et al., 2019; Craik et al., 2019; Al-Saegh et al., 2021; Sekkal et al., 2022; Yang et al., 2022). For example, some studies utilize convolutional neural network (CNN) to extract temporal and spatial features from EEG for different BCI tasks like motor imagery classification, emotion recognition and seizure detection (Schirrmeister et al., 2017; Sakhavi et al., 2018; Lawhern et al., 2018; Abdelhameed & Bayoumi, 2021; Ding et al., 2022). Long Short-Term Memory (LSTM) is also used for EEG feature extraction and classification on BCI tasks such as motor imagery classification and sleep staging (Wang et al., 2018; Phan et al., 2019). Some researchers propose CNN-LSTM to learn EEG features for motor imagery classification, emotion recognition and sleep staging (Supratak et al., 2017; Zhang et al., 2019; Dar et al., 2020; Li et al., 2022; Wang et al., 2023a), where CNN is usually used to extract local features and LSTM is utilized to capture global dependencies. Transformer architecture is also utilized to learn spatial-temporal feature for BCI tasks including emotion recognition, sleep staging and person identification (Song et al., 2021; Liu et al., 2021a; Du et al., 2022; Phan et al., 2022). To combine the strengths of CNN and transformer, some works devise a CNN-Transformer network for EEG classification (Song et al., 2022; Peh et al., 2022; Wang et al., 2023b; Zhou et al., 2024b; Wang et al., 2024c). In addition, some studies use graph neural networks learn spatial-temporal features from multi-channel EEG for various BCI tasks (Jia et al., 2020; 2021; Ding et al., 2023). These methods perform well on specific tasks or datasets, but collecting labeled EEG data is costly and labor-intensive. Meantime, the variations in EEG signal formats across different datasets pose challenges for enhancing model performance and generalizability.
@@
Brain Foundation Model. A foundation model (Bommasani et al., 2021) is a deep learning model trained on extensive data, typically using self-supervision at a large scale, and can be adapted, such as through fine-tuning, for various downstream tasks. Foundation models, such as BERT (Devlin et al., 2018), MAE (He et al., 2022), CLIP (Radford et al., 2021), GPT-4 (Achiam et al., 2023), SAM (Kirillov et al., 2023) and SORA (Brooks et al., 2024), have achieved remarkable success in computer vision, natural language, and multimodal, but there is still a vast unexplored potential for foundation models in brain signals. Some studies propose time series pre-trained models (Eldele et al., 2021; Zhang et al., 2022; Woo et al., 2024; Chen et al., 2024; Deng et al., 2024) based on contrastive learning or others for many scenarios (e.g., weather, traffic flow, exchange rates), one of which is brain signals. Some studies learn unsupervised representations from EEG by utilizing self-supervised learning pretext tasks to uncover the structure of clinical EEG signals (Banville et al., 2021). BENDER (Kostas et al., 2021) is proposed to address the problem of limited labeled data on EEG by using a contrastive self-supervised learning task to learn generic EEG representations. MAEEG (Chien et al., 2022) is masked auto-encoder for learning EEG representations by reconstruction-based self-supervised learning. BrainBERT (Wang et al., 2022) is a pre-training model for intracranial recordings, which learns a complex non-linear transformation of neural data by construction of the masked stereo-electroencephalographic (SEEG) spectrogram. Brant (Zhang et al., 2023) is a foundation model for intracranial neural signals which attends long-term dependency and captures spatial correlation across channels. Brant-2 (Yuan et al., 2024) represents an improved version of Brant, tailored for the integration of scalp EEG and intracranial EEG. Brant-X (Zhang et al., 2024) extends this framework to achieve alignment between different physiological signals and EEG. BIOT (Yang et al., 2023) is a generic biosignal learning model which enables joint pre-training and knowledge transfer across different biosignal datasets in the wild. EEG2Rep (Foumani et al., 2024) is a self-prediction approach for self-supervised representation learning from EEG, which predicting masked inputs in the latent representation space. LaBraM (Jiang et al., 2024) is a large brain model, which learns EEG generic representations by predicting the corresponding neural tokens of masked EEG patches. EEGPT (Wang et al., 2024a) a pretrained transformer model designed for universal EEG feature extraction, based on a mask-based dual self-supervised learning method for efficient feature extraction. However, most of existing EEG foundational model the spatial and temporal dependencies between all EEG patches together, ignoring the unique structural characteristics of EEG signals. Meanwhile, existing EEG foundation models have limited generalizability, performing well only on a limited range of downstream tasks.
@@ Appendix D More Details for Experimental Settings on Downstream BCI Tasks
D.1 More Details for Downstream Task
I. Emotion Recognition refers to detecting and interpreting emotional states or patterns based on EEG. We choose FACED (Chen et al., 2023) and SEED-V (Liu et al., 2021b) as downstream datasets for this task.
II. Motor Imagery Classification is a process that involves decoding or classifying motor imagery tasks based on brain activity patterns captured through EEG. We use PhysioNet-MI (Schalk et al., 2004) and SHU-MI (Goldberger et al., 2000b) to evaluate the performance of CBraMod on this task.
III. Sleep Staging, also known as sleep stage classification, categorizes different stages of sleep based on physiological signals (e.g. EEG) recorded during sleep. ISRUC (Khalighi et al., 2016) is a commonly used benchmark for sleep staging.
IV. Seizure Detection refers to the process of identifying and recognizing epileptic seizures based on physiological signals, such as EEG. Here, CHB-MIT (Shoeb, 2009) is choosed for this task.
V. Imagined Speech Classification involves the identification and interpretation of imagined speech or covert speech intentions based on brain activity patterns such as EEG. BCIC2020-3 (Jeong et al., 2022) is a dataset for this task on 2020 international brain–computer interface competition.
VI. Mental Disorder Diagnosis refers to the process of identifying and categorizing mental health states based on EEG. Mumtaz2016 (Mumtaz, 2016) is an EEG dataset collected from patients with major depressive disorder and normal controls.
VII. Vigilance Estimation refers to the assessment and measurement of an individual’s level of vigilance or sustained attention from EEG. SEED-VIG (Min et al., 2017) is dataset used to assess the level of vigilance.
VIII. Mental Stress Detection involves using EEG to identify an individual’s level of stress. MentalArithmetic (Zyma et al., 2019) is a dataset containing EEG recordings of subjects before and during the performance of mental arithmetic tasks.
IX. EEG Events Classification involves the categorization and identification of specific events or patterns in EEG. We choose TUEV (Obeid & Picone, 2016) for evaluation, which was consistent with BIOT (Yang et al., 2023) and LaBraM (Jiang et al., 2024).
X. Abnormal Detection refers to the process of identifying and detecting abnormal patterns or events in EEG. Similarly, TUAB (Obeid & Picone, 2016) is used for evaluation, consistent with BIOT (Yang et al., 2023) and LaBraM (Jiang et al., 2024).
D.2 Baselines
@@ E.9 Motor Imagery Classification
To further validate the performance of CBraMod on the motor imagery tasks, we conducted additional experiments on the BCIC-IV-2a (Brunner et al., 2008) dataset [^2] As the reviewer suggested, we have added an additional comparison experiment using the BCIC-IV-2a dataset. For the sake of narrative clarity, we did not include this dataset in the count of the 12 publicly available downstream datasets presented in the main text. Including this dataset, we have actually evaluated a total of 13 downstream datasets.. The BCI Competition IV Dataset 2a, provided by Graz University of Technology, comprises EEG recordings from 9 subjects performing 4 motor imagery tasks: imagining movements of the left hand, right hand, both feet, and tongue. Data were collected over two sessions on separate days using 22 Ag/AgCl electrodes at a sampling rate of 250 Hz. Each session consisted of 288 EEG trials, with 72 trials per task. We used [2, 6] seconds of each trial. A band-pass filtered (0.3 Hz–40 Hz) was applied to remove the low-frequency and high-frequency noise. We resample the EEG signals to 200 Hz and obtain 5,088 4-second samples. We use a strict subject-indenpendent train/validation/test strategy as other MI datasets in our paper. Subject 1-5, 6-7, 8-9 are used for training, validation, and test, respectively. The results are shown in Table 15. CBraMod continues to outperform existing methods on this dataset, further reinforcing the generalizability of our method.